OpenStudy (anonymous):

Check my answer! Determine if the expression is a perfect square and if it is, factor it. 16a^2 + 8a + 1 A. (a – 4)(a – 4) B. not a perfect square C. (4a + 1)2 D. (a + 4)(a – 4) My answer: B. You can read and check my reasoning below.

OpenStudy (anonymous):

\[16a^2 = 4a.~16a^2~is~a~perfect~\square.\]\[1=1^2. ~One~is~a~perfect~\square.\]

OpenStudy (anonymous):

b or d i think it is

OpenStudy (anonymous):

I believe 8a is supposed to be twice the product of the square roots of the first and last terms (AKA 16a^2 and 1).

OpenStudy (anonymous):

Oh, um, unless you have an explanation, please do not give out a direct answer. Thank you. @officialkeily

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

It's okay. :) Just remember to read the Code-of-Conduct for the rules. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Back to the problem, the expression would then be: \[8a = 2(4a \times 1)\]\[8a = (2 \times 4a)(2 \times 1)\]\[8a = 8a \times 2\]\[8a \neq 16a\]I guess that makes this expression not be a perfect square.

OpenStudy (anonymous):

@asnaseer This is my work so far. Is my progress correct?

DivineSolar (divinesolar):

I am pretty sure you are wrong :)

OpenStudy (anonymous):

:O :(

DivineSolar (divinesolar):

Factoring 16a2+8a+1 The first term is, 16a2 its coefficient is 16 . The middle term is, +8a its coefficient is 8 . The last term, "the constant", is +1 Step-1 : Multiply the coefficient of the first term by the constant 16 • 1 = 16 Step-2 : Find two factors of 16 whose sum equals the coefficient of the middle term, which is 8 . -16 + -1 = -17 -8 + -2 = -10 -4 + -4 = -8 -2 + -8 = -10 -1 + -16 = -17 1 + 16 = 17 2 + 8 = 10 4 + 4 = 8 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, 4 and 4 16a2 + 4a + 4a + 1 Step-4 : Add up the first 2 terms, pulling out like factors : 4a • (4a+1) Add up the last 2 terms, pulling out common factors : 1 • (4a+1) Step-5 : Add up the four terms of step 4 : (4a+1) • (4a+1) Which is the desired factorization

DivineSolar (divinesolar):

Does that help at all?

DivineSolar (divinesolar):

@BlossomCake ?

OpenStudy (anonymous):

Yes, it does help. But between steps 3 and 4, where does 16 go?

DivineSolar (divinesolar):

We no longer need it as you can see.

DivineSolar (divinesolar):

Keep in mind step 3 does say we are rewriting the expression :).

OpenStudy (anonymous):

Oh, yeah. Sorry. So which answer is correct then, D.?

OpenStudy (anonymous):

WAIT! Nevermind, the answer is C. Sorry that I am so slow. :')

OpenStudy (anonymous):

Thank you for all your help and explaining.

OpenStudy (anonymous):

*all of your

DivineSolar (divinesolar):

Sure thing.!

OpenStudy (anonymous):

:D